OpenStudy (anonymous):

If f(x) = xsquared + 1, find f(3). 10 7 4 2

OpenStudy (accessdenied):

just insert x=3 into f(x) f(x) = x^2 + 1 f(3) = 3^2 + 1 and evaluate